The sum of Stuart's and Tracy's age is 27. Stuart's age five year from now will be twice what Tracy's age was 2 years ago. How old is each person now?
Answer by stanbon(75887)

The sum of Stuart's and Tracy's age is 27. Stuart's age five year from now will be twice what Tracy's age was 2 years ago. How old is each person now?
----------
Equations:
s + t = 27 yrs.
s+5 = 2(t-2)
-----
s+t = 27
s-2t -9
-----------
Subtract and solve for "t":
3t = 36
t = 12 years (tracy's age now)
-----------
Solve for "s":
s + t = 27
s = 15 yrs
=================
Cheers,
Stan H.
================